(3xy+3y-4)dx+(x+1)^2dy=0


1
Expert's answer
2021-10-17T17:44:38-0400
"(3xy+3y-4)dx+(x+1)^2dy=0"

"(x+1)^2y'+3(x+1)y=4"

"y'+\dfrac{3}{x+1}y=\dfrac{4}{(x+1)^2}"

Integrating factor


"\mu(x)=e^{\int({3 \over x+1})dx}=(x+1)^3"

"(x+1)^3y'+3(x+1)^2y=4(x+1)"

"d((x+1)^3y)=4(x+1)dx"

Integrate


"\int d((x+1)^3y)=\int4(x+1)dx"

"(x+1)^3y=2(x+1)^2+C"

"y=\dfrac{2}{x+1}+\dfrac{C}{(x+1)^3}"